# Neural density functional theory of liquid-gas phase coexistence

### v1

@@ III.1 Bulk pair correlation functions
Figure 2: Bulk pair correlation functions for constant supercritical temperature $k_{B}T/\varepsilon=1.5$ and different bulk densities $\rho_{b}\sigma^{3}=0.1,0.2,0.3,0.4,0.5,0.6$, and $0.7$. Shown are (a) the planar two-body direct correlation function $\bar{c}_{2}^{b}(x)$ along with the radial representation $c_{2}^{b}(r)$ (inset), (b) its Fourier transform $\tilde{c}_{2}^{b}(k)$ and the total correlation function $\tilde{h}(k)$ in Fourier space (inset) as obtained from the OZ equation (13), (c) the static structure factor $S(k)$ and the radial distribution function $g(r)$ (inset).
@@
The total correlation function is obtained via the Ornstein-Zernike (OZ) route. A one-dimensional Fourier transform of $\bar{c}_{2}(x)$ yields the radial quantity $\tilde{c}_{2}^{b}(k)$ in Fourier space (indicated by the tilde). The Ornstein-Zernike equation
@@
Results for the various bulk pair correlation functions are shown in Fig. 2 for a supercritical temperature of $k_{B}T/\varepsilon=1.5$ and different bulk densities. See also Ref. [28] for comparative data obtained with isothermal training and from other methods, e.g. pair-correlation matching [23], and Ref. [25] for pair correlations of the hard sphere fluid. From Fig. 2 we note: i) noisy artifacts arise in $c_{2}^{b}(r)$ and $g(r)$ for $r<\sigma$ due to numerical intricacies associated with the transformations (12) and (15), ii) $c_{2}^{b}(r)$ and $\bar{c}_{2}^{b}(x)$ are independent of density $\rho_{b}$ for $r>\sigma$, reflecting the fact that $c_{2}^{b}(r)$ quickly reaches its asymptotic limit $-\beta\phi(r)$ [65], and iii) $S(k)$ exceeds $2.0$ for small wave numbers $k$ at reduced densities $\rho_{b}\sigma^{3}=0.2,0.3$, a possible sign of the approach to a critical point.
@@ III.2 Lines of maximal isothermal compressibility and correlation length, Fisher-Widom line and spinodal
We turn now to the asymptotic (large $r$) decay of $h(r)$. Specifically, we determine both the Widom line of maximal true correlation length, middle column, and the Fisher-Widom crossover line, third column of Fig. 3. A pole analysis of the total correlation function determines the asymptotic behaviour of $h(r)$ at long range, $r\rightarrow\infty$, where either (damped) oscillatory or monotonic decay is exhibited [65, 66, 67]. In Fourier space, the poles $\alpha=\alpha_{1}+i\alpha_{0}$ of the total pair correlation function are readily determined by the OZ equation (13), leading back to the analysis of the direct correlation function $\tilde{c}_{2}^{b}(k)$. The zeroes of the denominator $1-\rho_{b}\tilde{c}_{2}^{b}(\alpha)$ in Eq. (13) yield the conditions [65]
@@
which is plotted for a range of bulk densities and temperatures in the second column in Fig. 3. The behaviour of $\xi$ is consistent with the results for the isothermal compressibility $\chi_{T}$: the correlation length increases rapidly when approaching the critical region and indicates a spinodal where $\xi$ diverges when continuing to lower the temperature. Tracing $\xi\rightarrow\infty$ or $\chi_{T}\rightarrow\infty$ yields numerically identical spinodals. Of course, this is expected since setting $\alpha_{0}^{\mathrm{mon}}$ to zero in Eq. (21) is equivalent to requiring the right hand side of Eq. (16) to vanish. Within the spinodal, Eq. (21) yields no solution and $\xi$ is undefined.
@@ III.3 Liquid-gas coexistence, binodal and estimate of critical point
One can attempt to fit the binodal and determine the critical point in various ways. Following Wilding [57], we exclude the near critical region and take the neural coexistence densities within the temperature range $0.95\leq k_{B}T/\varepsilon\leq 1.15$ to fit the binodal via [57]
@@
Unlike in Ref. [57], the critical point is, a priori, undetermined in Eq. (23); $\rho_{c}$ and $T_{c}^{*}$ are to be deduced in the fit along with all other parameters. To demonstrate the robustness of the fitting procedure, we also keep the (critical) exponent $\beta$ as a free parameter, although we bear in mind the Ising result $\beta=0.32630(22)$ [68]. The results of the binodal fit are shown in Fig. 4 and agree very well with the data from highly accurate simulations [57]. This procedure, which focuses on coexistence densities, attempts to avoid some of the intricacies of the critical region that are also pertinent for the neural functional (see Sec. III.5). We note that the resulting value of the ‘critical’ exponent $\beta$ is, from full training, 0.330 which lies close to the Ising value [68], demonstrating that the fit is meaningful as well as indicating possible beyond-mean-field character of the neural functional. Note also that the corresponding parameters/results from Wilding’s MC simulations and fitting [57]: $k_{B}T_{c}/\varepsilon=1.188$, $\rho_{c}\sigma^{3}=0.320$, $a\sigma^{3}=0.182$, $b\sigma^{3}=0.523$, $\beta=0.326$, are close to the present.
Following the work of Panagiotopoulos [58], an alternative estimate of the critical temperature $T_{c}$ and density $\rho_{c}$ proceeds by a regression and extrapolation of the cubed difference of coexistence densities and of the rectilinear diameter law, respectively. As before, coexistence data in the close vicinity of the critical point needs to be excluded. The thrust of this approach is the prediction that $\rho_{l}-\rho_{g}$ should decay approximately as $(T_{c}-T)^{1/3}$ near the critical point rather than as $(T_{c}-T)^{1/2}$, which is the mean-field prediction. Fig. 5(a) indicates that our neural functional yields results consistent with non-mean-field behaviour and the estimate of the critical temperature is close to the simulation result. Fig. 5(b) plots the rectilinear diameter versus temperature which allows for an estimate of the critical density that is again close to the simulation result. The results shown in Fig. 5 are consistent with the previous fit to the binodal using Eq. (23).
@@ III.4 Bulk equation of state, liquid-gas interface and surface tension
Figure 6: Bulk equation of state: pressure $P$ as a function of the bulk density $\rho_{b}$ for temperatures $0.7\leq k_{B}T/\varepsilon\leq 1.3$. The inset shows the same data plotted with respect to the inverse bulk density, i.e. proportional to the system volume. Note the appearance of a van der Waals loop at subcritical temperatures.
Figure 7: (a) Liquid-gas interfacial density profiles as obtained from solution of Eq. (8) for a phase-separated system upon keeping the mean density fixed. Results are shown for five temperatures $k_{B}T/\varepsilon=0.7,0.8,0.9,1.0$ and $1.1$. The inset indicates the respective bulk coexistence densities, the fitted binodal (dotted black line) and critical point (black dot) reproduced from Fig. 4; the horizontal dotted pink line shows the temperature cutoff of $k_{B}T/\varepsilon=1.0$ during training. Oscillations of the interfacial density profile that extend into the liquid domain are visible for the two lowest temperatures. (b) Surface tension $\gamma_{lg}$ of the liquid-gas interface as determined from functional line integration of interfacial density profiles. Simulation results are taken from Refs. [61, 60, 62]. The cutoff for training, $k_{B}T/\varepsilon=1.0$, is indicated again, now by the vertical dotted pink line.
@@
where $c_{1}^{b}=c_{1}(x;[\rho_{b}],T)$ follows from direct evaluation of the neural correlation functional with constant density input and at arbitrary position $x$ due to translational invariance. Note that the excess chemical potential can be identified as $\mu_{\mathrm{exc}}=-k_{B}Tc_{1}^{b}$. We show the neural prediction for the equation of state $P(\rho_{b},T)$ for a range of temperatures in Fig. 6. Strikingly, a van der Waals loop emerges for subcritical temperatures, i.e. for $k_{B}T/\varepsilon<1.2$. Note that negative pressures arise at low temperatures. The inset to Fig. 6 plots pressure versus volume. We have checked that a Maxwell equal area construction confirms the values of the coexisting liquid and gas densities as determined in Sec. III.3 within the expected numerical accuracy.
@@
with system volume $V$, lateral system area $A$ and pressure $P$ at coexistence as given by Eq. (24). We note that subtle discrepancies in the values of the liquid and gas pressures $P_{l}$ and $P_{g}$ arise for low temperatures due to the accumulation of numerical errors when integrating through the van der Waals loop in Eq. (24). To alleviate this numerical issue, we compute $PV=P_{l}V_{l}+P_{g}V_{g}$ and employ this relation consistently in Eq. (25); $V_{l}$ and $V_{g}$ denote the volumes of liquid and gas domains, respectively. This procedure is inline with the definition of the surface tension $\gamma_{lg}$ as a genuine excess quantity. We also note that the pressures predicted at bulk coexistence are generally small, with values in the order of $10^{-2}\varepsilon/\sigma^{3}$.
@@ III.5 Critical region and Ornstein-Zernike plots
Figure 8: OZ plot of the inverse structure factor $S$ with respect to $k^{2}$ for supercritical temperatures at constant critical density $\rho_{c}$, from which one can determine $S(0)$, the short-range correlation length $R$, and the OZ correlation length $\xi_{\mathrm{OZ}}$—see text. Plots of $S(0)$ and $\xi_{\mathrm{OZ}}$ with respect to $|T-T_{c}|/T_{c}$ determine the critical exponents $\gamma$ and $\nu$ (inset); our neural functional yields the mean field values as shown. The true correlation length $\xi$, calculated from Eq. (22), lies very close to $\xi_{\mathrm{OZ}}$ for temperatures close to $T_{c}$, as shown (in green) in the inset.
@@
Although the neural functional is straightforward to evaluate for state points in the vicinity of the critical point, the resulting predictions must be treated with particular caution. The limitations are already apparent by comparing results obtained for the spinodal, see the isothermal compressibility $\chi_{T}$ and correlation length $\xi$ in Fig. 3, with the estimate of the critical point from the binodal fit in Fig. 4. One observes that the critical point obtained from the latter does not match exactly the top of the spinodal from the former. To be precise, we get $k_{B}T_{c}/\varepsilon\approx 1.188$ from the binodal fit in Fig. 4 but $k_{B}T_{c}/\varepsilon\approx 1.203$ from the locus of either $\chi_{T}^{-1}=0$ or $\xi^{-1}=0$ in Fig. 3. We should first place these results in perspective: we are focusing on estimates of the critical point, obtained via our neural functional that had no prior knowledge of the existence of phase separation let alone a critical point, which lie within about one percent of the most accurate, independent, simulation study of the critical point of the LJ fluid [57] which finds $k_{B}T_{c}/\varepsilon=1.188$.
@@
In the following, we investigate the subtle issues that arise when inferring properties in the critical region. Recall that the neural functional yields coexisting densities that are fitted better, within our procedures, by a non mean-field critical exponent close to the Ising value $\beta\approx 0.33$ than by the mean-field result $\beta=0.5$, see Sec. III.3. However, the physics is more complex. For example, Panagiotopoulos [58] argues that certain properties in the critical region could be forced to appear mean-field-like because of the finite size of the simulation. The basic idea is that one cannot access correlation lengths that are larger than or of the same order as the simulation box size. In our study we provide training data from MC simulations, of planar density profiles in boxes of lateral size $L<20\sigma$. It is clear that long correlation lengths will be suppressed, and one might therefore expect the neural functional to inherit mean-field critical behaviour. We choose to examine this issue by performing an OZ plot of the static structure factor, cf. Sec. III.1 and Eq. (14). Specifically we employ the traditional OZ description $S(k)=S(0)/(1+\xi_{\mathrm{OZ}}^{2}k^{2})$ for small wavenumbers $k$. The inverse of $S(k)$ is plotted against $k^{2}$, at supercritical temperatures approaching $T_{c}$, and at fixed critical density $\rho_{c}\sigma^{3}=0.322$, in Fig. 8. Note that the reference value of the critical temperature refers here to $k_{B}T_{c}/\varepsilon\approx 1.2031$ as given by the maximum of the spinodal, defined, of course, by $01/S(0)=$ 0.
@@
Defining the critical exponent $\nu$ by $\xi_{\mathrm{OZ}}\sim|T-T_{c}|^{-\nu}$ it follows that our analysis predicts $\gamma=2\nu$, since $R$ remains finite. This is implies a (Fisher) exponent $\eta=0$. The inset in Fig. 8 shows a log-log plot of $S(0)$ and the OZ correlation length $\xi_{\mathrm{OZ}}$ as functions of the reduced supercritical temperature difference $(T-T_{c})/T_{c}$. Our neural functional results yield the mean-field critical exponents $\gamma=1$ and $\nu=0.5$ to high accuracy, in keeping with ideas of Panagiotopoulos [58] and others.
We also investigated the critical behaviour of the true correlation length $\xi$ as given by the pole analysis, see Sec. III.2 and Eq. (22). As expected, see the inset, $\xi$ and $\xi_{\mathrm{OZ}}$ are almost identical near the critical point. Hence, we find the true correlation length also exhibits mean-field behaviour with critical exponent $\nu=0.5$. That these two different correlation lengths obtained from the neural functional are consistent in the critical region, albeit both exhibiting mean-field scaling, gives us confidence in our numerical implementations.

### v2

@@ III.1 Bulk pair correlation functions
Figure 2: Bulk pair correlation functions from the neural functional for constant supercritical temperature $k_{B}T/\varepsilon=1.5$ and different bulk densities $\rho_{b}\sigma^{3}=0.1,0.2,0.3,0.4,0.5,0.6$, and $0.7$, as indicated by ticks on the colorscale and labeling of the curves. Shown are (a) the planar two-body direct correlation function $\bar{c}_{2}^{b}(x)$ along with the radial representation $c_{2}^{b}(r)$ (inset), (b) its Fourier transform $\tilde{c}_{2}^{b}(k)$ and the total correlation function $\tilde{h}(k)$ in Fourier space (inset) as obtained from the OZ equation (13), (c) the static structure factor $S(k)$ and the radial distribution function $g(r)$ (inset). For three bulk densities $\rho_{b}\sigma^{3}=0.1,0.4$, and $0.7$, we show results for each quantity (dotted lines), obtained from Fourier transform and OZ inversion of bulk grand canonical MC simulation data for $g(r)$.
@@
The total correlation function is obtained via the Ornstein-Zernike (OZ) route. A one-dimensional Fourier transform of $\bar{c}_{2}^{b}(x)$ yields the radial quantity $\tilde{c}_{2}^{b}(k)$ in Fourier space (indicated by the tilde). The Ornstein-Zernike equation
@@
Results for the various bulk pair correlation functions are shown in Fig. 2 for a supercritical temperature of $k_{B}T/\varepsilon=1.5$ and different bulk densities. See also Ref. [28] for comparative data obtained with isothermal training and from other methods, e.g. pair-correlation matching [23], and Ref. [25] for pair correlations of the hard sphere fluid. From Fig. 2 we note: i) noisy artifacts arise in $c_{2}^{b}(r)$ and $g(r)$ for $r<\sigma$ due to numerical intricacies associated with the transformations (12) and (15), ii) $c_{2}^{b}(r)$ and $\bar{c}_{2}^{b}(x)$ are independent of density $\rho_{b}$ for $r>\sigma$, reflecting the fact that $c_{2}^{b}(r)$ quickly reaches its asymptotic limit $-\beta\phi(r)$ [67], iii) $S(k)$ exceeds $2.0$ for small wave numbers $k$ at reduced densities $\rho_{b}\sigma^{3}=0.2,0.3$, a possible sign of the approach to a critical point, and iv) the neural predictions match very closely the simulation results extracted from separate grand canonical MC for $g(r)$—see dotted lines referring to three specific densities in Fig. 2. We emphasize that no information regarding pair correlations was incorporated during training.
@@ III.2 Lines of maximal isothermal compressibility and correlation length, Fisher-Widom line and spinodal
We turn now to the asymptotic (large $r$) decay of $h(r)$. Specifically, we determine both the Widom line of maximal true correlation length, middle column, and the Fisher-Widom crossover line, third column of Fig. 3. A pole analysis of the total correlation function determines the asymptotic behavior of $h(r)$ at long range, $r\rightarrow\infty$, where either (damped) oscillatory or monotonic decay is exhibited [67, 68, 69]. In Fourier space, the poles $\alpha=\alpha_{1}+i\alpha_{0}$ of the total pair correlation function are readily determined by the OZ equation (13), leading back to the analysis of the direct correlation function $\tilde{c}_{2}^{b}(k)$. The zeroes of the denominator $1-\rho_{b}\tilde{c}_{2}^{b}(\alpha)$ in Eq. (13) yield the conditions [67]
@@
which is plotted for a range of bulk densities and temperatures in the second column in Fig. 3. The behavior of $\xi$ is consistent with the results for the isothermal compressibility $\chi_{T}$: the correlation length increases rapidly when approaching the critical region and indicates a spinodal where $\xi$ diverges when continuing to lower the temperature. Tracing $\xi\rightarrow\infty$ or $\chi_{T}\rightarrow\infty$ yields numerically identical spinodals. Of course, this is expected since setting $\alpha_{0}^{\mathrm{mon}}$ to zero in Eq. (21) is equivalent to requiring the right hand side of Eq. (16) to vanish. Within the spinodal, Eq. (21) yields no solution and $\xi$ is undefined.
@@ III.3 Liquid-gas coexistence, binodal and estimate of critical point
One can attempt to fit the binodal and determine the critical point in various ways that can incorporate critical exponents beyond mean-field. Following Wilding [59], we exclude the near critical region and take the neural coexistence densities within the temperature range $0.95\leq k_{B}T/\varepsilon\leq 1.15$ to fit the binodal via [59]
@@
Unlike in Ref. [59], the critical point is, a priori, undetermined in Eq. (23); $\rho_{c}$ and $T_{c}^{*}$ are to be deduced in the fit along with all other parameters. To demonstrate the robustness of the fitting procedure, we also keep the (critical) exponent $\beta$ as a free parameter, although we bear in mind the Ising result $\beta=0.32630(22)$ [70]. The results of the binodal fit are shown in Fig. 4 and agree very well with the data from highly accurate simulations [59]. This procedure, which focuses on coexistence densities at temperatures sufficiently far below the critical point such that the correlation length $\xi$ is less than the system size, attempts to avoid some of the intricacies of the critical region that are also pertinent for the neural functional (see Sec. III.5). We note that the resulting value of the ‘critical’ exponent $\beta$ is, from full training, 0.330 which lies close to the Ising value [70], demonstrating that the fit is meaningful as well as indicating possible beyond-mean-field character of the neural functional. Note also that the corresponding parameters/results from Wilding’s MC simulations and fitting [59]: $k_{B}T_{c}/\varepsilon=1.188$, $\rho_{c}\sigma^{3}=0.320$, $a\sigma^{3}=0.182$, $b\sigma^{3}=0.523$, $\beta=0.326$, are close to the present.
Following the work of Panagiotopoulos [60], an alternative estimate of the critical temperature $T_{c}$ and density $\rho_{c}$ proceeds by a regression and extrapolation of the cubed difference of coexistence densities and of the rectilinear diameter law, respectively. As before, coexistence data in the close vicinity of the critical point needs to be excluded. The thrust of this approach is the prediction that $\rho_{l}-\rho_{g}$ should decay approximately as $(T_{c}-T)^{1/3}$ near the critical point rather than as $(T_{c}-T)^{1/2}$, which is the mean-field prediction. Fig. 5(a) indicates that our neural functional yields results consistent with non-mean-field behavior and the estimate of the critical temperature is close to the simulation result. Fig. 5(b) plots the rectilinear diameter versus temperature which allows for an estimate of the critical density that is again close to the simulation result. The results shown in Fig. 5 are consistent with the previous fit to the binodal using Eq. (23).
@@ III.4 Bulk equation of state, liquid-gas interface and surface tension
Figure 6: Bulk equation of state: pressure $P$ as a function of the bulk density $\rho_{b}$ for temperatures $k_{B}T/\varepsilon=0.7,0.8,0.9,1.0,1.1,1.2$, and $1.3$, as labeled and indicated on the colorscale. The inset shows the same data plotted with respect to the inverse bulk density, i.e. proportional to the system volume. Note the appearance of a van der Waals loop at subcritical temperatures.
Figure 7: (a) Liquid-gas interfacial density profiles as obtained from solution of Eq. (8) for a phase-separated system upon keeping the mean density fixed. Results are shown for five temperatures $k_{B}T/\varepsilon=0.7,0.8,0.9,1.0$, and $1.1$, see labels. The inset indicates the respective bulk coexistence densities, the fitted binodal (dotted black line) and critical point (black dot) reproduced from Fig. 4; the horizontal dotted pink line shows the temperature cutoff of $k_{B}T/\varepsilon=1.0$ during training. Oscillations of the interfacial density profile that extend into the liquid domain are visible for the two lowest temperatures. (b) Surface tension $\gamma_{lg}$ of the liquid-gas interface as determined from functional line integration of interfacial density profiles. Simulation results are taken from Refs. [63, 62, 64]. The cutoff for training, $k_{B}T/\varepsilon=1.0$, is indicated again, now by the vertical dotted pink line. Fitting to neural data yields the effective exponent for the tension $\tilde{\mu}\approx 1.26$—see text.
@@
where $c_{1}^{b}=c_{1}(x;[\rho_{b}],T)$ follows from direct evaluation of the neural correlation functional with constant density input and at arbitrary position $x$ due to translational invariance. Note that the excess chemical potential can be identified as $\mu_{\mathrm{exc}}=-k_{B}Tc_{1}^{b}$. We show the neural prediction for the equation of state $P(\rho_{b},T)$, obtained from Eq. (24), for a range of temperatures in Fig. 6. Note that a van der Waals loop emerges for subcritical temperatures, i.e. for $k_{B}T/\varepsilon\lesssim 1.2$ and that negative pressures arise at low temperatures. The inset to Fig. 6 plots pressure versus volume. Performing a Maxwell equal area construction yields, within the expected numerical accuracy, values of the coexisting liquid and gas densities equal to those determined in Sec. III.3 by direct solution of the Euler-Lagrange equation for an inhomogeneous phase-separated system. This attests to the consistency of the neural functional.
@@
with system volume $V$, lateral system area $A$ and pressure $P$ at coexistence as given by Eq. (24). We note that subtle discrepancies in the values of the liquid and gas pressures $P_{l}$ and $P_{g}$ arise for low temperatures due to the accumulation of numerical errors when integrating through the van der Waals loop in Eq. (24). To alleviate this numerical issue, we compute $PV=P_{l}V_{l}+P_{g}V_{g}$ and employ this relation consistently in Eq. (25); $V_{l}$ and $V_{g}$ denote the volumes of liquid and gas domains, respectively. This procedure is inline with the definition of the surface tension $\gamma_{lg}$ as a genuine excess quantity. We also note that the pressures predicted at bulk coexistence are generally small, with values of the order of $10^{-2}\varepsilon/\sigma^{3}$.
@@
The data in Fig. 7 obtained for a range of subcritical temperatures allows us to investigate the near-critical behavior of the surface tension. Recall that the latter should vanish as $\gamma_{lg}\sim(T_{c}-T)^{\tilde{\mu}}$ with a critical exponent $\tilde{\mu}$, given by $\tilde{\mu}=2\nu\approx 1.26$ for the three-dimensional Ising case, where the correlation length diverges as $\xi\sim(T_{c}-T)^{-\nu}$. We proceed analogously to Sec. III.3 and perform a fit of subcritical data within the same temperature range $0.95\leq k_{B}T/\varepsilon\leq 1.15$. Importantly, we thereby fix the critical temperature $k_{B}T_{c}/\varepsilon=1.1883$ as obtained from the binodal regression in Fig. 4. This choice of fitting procedure yields the ‘effective’ critical exponent $\tilde{\mu}\approx 1.26$, consistent with the Ising value. Recall that the mean-field exponent is $\tilde{\mu}=3/2$. We also examined the ‘10-90’ width of the liquid-gas interface in the same temperature range. This appears to diverge in the same fashion as the correlation length, i.e. with exponent $\nu\approx 0.63$, as expected from scaling arguments.
We conclude that using density profiles of the free interface, in a suitably chosen range of subcritical temperatures, allows one to perform fits to both the coexistence densities, cf. Fig. 4, and the surface tension, cf. Fig. 7, that suggest non-mean-field behavior. Of course, in making such fits one is deliberately avoiding the direct evaluation of the neural functional for temperatures very close to $T_{c}$, where the correlation length becomes very long. We might expect results to cross over to mean-field behavior, as a result of the finite size of our systems as we lay out in the following section.
@@ III.5 Critical region and Ornstein-Zernike plots
Figure 8: OZ plot of the inverse structure factor $S$ with respect to $k^{2}$ for supercritical temperatures at constant critical density $\rho_{c}$, from which one can determine $S(0)$, the short-range correlation length $R$, and the OZ correlation length $\xi_{\mathrm{OZ}}$—see text. Plots of $S(0)$ and $\xi_{\mathrm{OZ}}$ with respect to $|T-T_{c}|/T_{c}$ determine the critical exponents $\gamma$ and $\nu$ (inset); our neural functional yields the mean-field values as shown. The true correlation length $\xi$, calculated from Eq. (22), lies very close to $\xi_{\mathrm{OZ}}$ for temperatures close to $T_{c}$, as shown (in green) in the inset.
@@
Although the neural functional is straightforward to evaluate for state points in the vicinity of the critical point, the resulting predictions must be treated with particular caution. The limitations are already apparent by comparing results obtained for the spinodal, see the isothermal compressibility $\chi_{T}$ and correlation length $\xi$ in Fig. 3, and the estimate of the critical point from fitting to the binodal using Eq. (23) in Fig. 4. One observes that the critical point obtained from the latter does not match exactly the top of the spinodal from the former. To be precise, we get $k_{B}T_{c}/\varepsilon\approx 1.188$ from the binodal fit in Fig. 4 but $k_{B}T_{c}/\varepsilon\approx 1.203$ from the locus of either $\chi_{T}^{-1}=0$ or $\xi^{-1}=0$ in Fig. 3. It is important to note that simply tracing the neural binodal (blue circles) in Fig. 4 will also yield the identical critical temperature $k_{B}T_{c}/\varepsilon\approx 1.203$ and critical density as obtained from the spinodal, as is expected from a DFT approach. On the other hand, our determination of the critical temperature, $k_{B}T_{c}/\varepsilon\approx 1.188$, obtained by fitting to a suitably chosen range of subcritical data, is inspired by simulation studies [60, 59] and deliberately invokes additional physics by excluding results in the near-critical region that will be beset by finite-size artifacts.
We first should place our results in perspective: we are focusing on estimates of the critical point, obtained via our neural functional that had no prior knowledge of the existence of phase separation let alone a critical point, which lie within about one percent of the most accurate, independent, simulation study of the critical point of the LJ fluid [59] which finds $k_{B}T_{c}/\varepsilon=1.188$.
@@
In the following, we scrutinize further issues that arise when inferring properties in the critical region. Recall that, for a suitably chosen temperature range, the neural functional yields coexisting densities that are fitted better, within our procedures, by a non mean-field critical exponent close to the Ising value $\beta\approx 0.33$ than by the mean-field result $\beta=1/2$, see Sec. III.3. Similarly, for the surface tension we find via an analogous procedure the critical exponent $\tilde{\mu}\approx 1.26$ instead of the mean-field prediction $\tilde{\mu}=3/2$. Of course, when approaching $T_{c}$ very closely the physics is more subtle. For example, Panagiotopoulos [60] argues that certain properties in the critical region could be forced to appear mean-field-like because of the finite size of the simulation. The basic idea is that one cannot access correlation lengths that are larger than or of the same order as the simulation box size. In our study we provide training data from MC simulations of planar density profiles in boxes of lateral size $L<20\sigma$. It is clear that large correlation lengths will be suppressed, and one might therefore expect the neural functional to inherit mean-field critical behavior.
We choose to examine this issue by performing an OZ plot of the static structure factor, cf. Sec. III.1 and Eq. (14), in the super-critical regime, employing the traditional OZ description $S(k)=S(0)/(1+\xi_{\mathrm{OZ}}^{2}k^{2})$ for small wavenumbers $k$. The inverse of $S(k)$ is plotted against $k^{2}$, at supercritical temperatures approaching $T_{c}$, and at fixed critical density $\rho_{c}\sigma^{3}=0.322$, in Fig. 8. Note that the reference value of the critical temperature refers here to $k_{B}T_{c}/\varepsilon\approx 1.2031$ as given by the maximum of the spinodal, defined, of course, by $01/S(0)=$ 0.
@@
Defining the critical exponent $\nu$ by $\xi_{\mathrm{OZ}}\sim|T-T_{c}|^{-\nu}$ it follows that our analysis predicts $\gamma=2\nu$, since $R$ remains finite. This implies a (Fisher) exponent $\eta=0$. The inset in Fig. 8 shows a log-log plot of $S(0)$ and the OZ correlation length $\xi_{\mathrm{OZ}}$ as functions of the reduced supercritical temperature difference $(T-T_{c})/T_{c}$. Our neural functional results yield the mean-field critical exponents $\gamma=1$ and $\nu=0.5$ to high accuracy, in keeping with ideas of Panagiotopoulos [60] and others.
We also investigated the critical behavior of the true correlation length $\xi$ as given by the pole analysis, see Sec. III.2 and Eq. (22). As expected, see the inset, $\xi$ and $\xi_{\mathrm{OZ}}$ are almost identical near the critical point. Hence, we find the true correlation length also exhibits mean-field behavior with critical exponent $\nu=0.5$. That these two different correlation lengths obtained from the neural functional are consistent in the critical region, albeit both exhibiting mean-field scaling, gives us confidence in our numerical implementations.